# Non-Homophilic Graph Pre-Training and Prompt Learning

### v4

@@ Abstract.
Graphs are ubiquitous for modeling complex relationships between objects across various fields. Graph neural networks (GNNs) have become a mainstream technique for graph-based applications, but their performance heavily relies on abundant labeled data. To reduce labeling requirement, pre-training and prompt learning has become a popular alternative. However, most existing prompt methods do not differentiate homophilic and heterophilic characteristics of real-world graphs. In particular, many real-world graphs are non-homophilic, not strictly or uniformly homophilic with mixing homophilic and heterophilic patterns, exhibiting varying non-homophilic characteristics across graphs and nodes. In this paper, we propose ProNoG, a novel pre-training and prompt learning framework for such non-homophilic graphs. First, we analyze existing graph pre-training methods, providing theoretical insights into the choice of pre-training tasks. Second, recognizing that each node exhibits unique non-homophilic characteristics, we propose a conditional network to characterize the node-specific patterns in downstream tasks. Finally, we thoroughly evaluate and analyze ProNoG through extensive experiments on ten public datasets. Codes are available at https://github.com/Jaygagaga/ProNoG/tree/main.
Graph mining, non-homophilic graph, prompt learning, pre-training, few-shot learning.
Co-first authors. Work was done while at the University of Tokyo.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Web miningccs: Information systems Data miningccs: Computing methodologies Learning latent representations
@@ 1. Introduction
Graph data are pervasive in real-world applications, such as citation networks (Kanakia et al., 2019; Xiong et al., 2017), social networks (Ji et al., 2023; Zhang et al., 2023), and molecular graphs (Wang et al., 2023; Lee et al., 2023). Traditional methods typically train graph neural networks (GNNs) (Kipf and Welling, 2017; Veličković et al., 2018) or graph transformers (Ying et al., 2021; Yun et al., 2019) in a supervised manner. However, they require re-training and substantial labeled data for each specific task.
@@
To mitigate the limitations of supervised methods, pre-training methods have gained significant traction (Hu et al., 2020a; Velickovic et al., 2019; You et al., 2020; Qiu et al., 2020). They first learn universal, task-independent properties from unlabeled graphs, and then fine-tune the pre-trained models to various downstream tasks using task-specific labels (Velickovic et al., 2019; You et al., 2020). However, a significant gap occurs between the pre-training objectives and downstream tasks, resulting in suboptimal performance (Yu et al., 2024a; Sun et al., 2023b). Moreover, fine-tuning large pre-trained models is costly and still requires sufficient task-specific labels to prevent overfitting. As an alternative to fine-tuning, prompt learning has emerged as a popular parameter-efficient technique for adaptation to downstream tasks (Sun et al., 2022; Liu et al., 2023a; Sun et al., 2023a; Tan et al., 2023; Fang et al., 2024; Yu et al., 2024c). They first utilize a universal template to unify pre-training and downstream tasks. Then, a learnable prompt is employed to modify the input features or hidden embeddings of the pre-trained model to align with the downstream task without updating the pre-trained weights. Since a prompt has far fewer parameters than the pre-trained model, prompt learning can be especially effective in low-resource settings (Yu et al., 2024a).
@@
However, current graph “pre-train, prompt” approaches rely on the homophily assumption or overlook the presence of heterophilic edges. Specifically, the homophily assumption (Ma et al., 2022; Zhu et al., 2020b) states that neighboring nodes should share the same labels, whereas heterophily refers to the opposite scenario where two neighboring nodes have different labels. We observe that real-world graphs are typically non-homophilic, meaning they are neither strictly or uniformly homophilic and mix both homophilic and heterophilic patterns (Xiao et al., 2022, 2023). In this work, we investigate the pre-training and prompt learning methodology for non-homophilic graphs. We first revisit existing graph pre-training methods for such graphs, followed by proposing a Prompt learning framework for Non-homophilic Graphs (or ProNoG in short). The solution is non-trivial, as the notion of homophily encompasses two key aspects, each with its own unique challenge.
@@
First, different graphs exhibit varying degrees of non-homophily. As shown in Fig. 1(a), the Cora citation network that is generally considered largely homophilic with 81% homophilic edges [^1] Defined as edges connecting two nodes of the same label; see Eq. (1) in Sect. 3., whereas the Wisconsin webpage graph links different kinds of webpages, which is highly heterophilic with only 21% homophilic edges. Moreover, the non-homophilic characteristics of a graph also depends on the target label. For example, in a dating network shown in Fig. 1(a), taking gender as the node label, the graph is more heterophilic with $2/7$ homophilic edges. However, taking hobbies as the node label, the graph becomes more homophilic with $4/7$ homophilic edges. Hence, how do we pre-train a graph model irrespective of the graph’s homophily characteristics? In this work, we propose definitions for homophily tasks and homophily samples. We show that pre-training with non-homophily samples increases the loss of any homophily task. Meanwhile, a less homophilic graph results in a higher number of non-homophily samples, subsequently increasing the pre-training loss for homophily tasks. This motivates us to move away from homophily tasks for graph pre-training (Sun et al., 2022; Liu et al., 2023a) and instead choose a non-homophily task (You et al., 2020; Xiao et al., 2022).
@@
Second, different nodes within the same graph are distributed differently in terms of their non-homophilic characteristics. As shown in Fig. 1(c), on both Cora and Cornell, their nodes have a diverse homophily ratios [^2] Defined as the fraction of a node’s neighbors with the same label; see Eq. (2) in Sect. 3.. Hence, how do we capture the fine-grained, node-specific non-homophilic characteristics? Due to the diverse characteristics across nodes, a one-size-fits-all solution for all nodes would be inadequate. However, existing approaches generally apply a single prompt to all nodes (Sun et al., 2022; Liu et al., 2023a; Sun et al., 2023a; Fang et al., 2024), treating all nodes uniformly. Thus, these methods overlook the fine-grained node-wise non-homophilic characteristics, leading to suboptimal performance. For example, a standard graph prompt learning approach (Liu et al., 2023a) generally performs worse when the homophily ratios of nodes decrease, as shown in Fig. 1(d),even with a non-homophily pretext task (You et al., 2020). Though some recent works (Tan et al., 2023; Chen et al., 2023) have proposed node-specific prompts, they are not designed to account for the variation in nodes’ non-homophilic charactersitics. Inspired by conditional prompt learning (Zhou et al., 2022a), we propose generating a unique prompt from each node with a conditional network (condition-net) to capture the fine-grained, distinct characteristics of each node. We first capture the non-homophilic patterns of each node by reading out its multi-hop neighborhood. Then, conditioned on these non-homophilic patterns, the condition-net produces a series of prompts, one for each node that reflects its varying non-homophilic characteristics. These prompts can adjust the node embeddings to better align them with the downstream task.
In summary, the contributions of this work are threefold: (1) We observe varying degrees of homophily across graphs, which motivates us to revisit graph pre-training tasks. We provided theoretical insights which guide us to choose non-homophily tasks for graph pre-training. (2) We further observe that, within the same graph, different nodes have diverse distributions of non-homophilic characteristics. To adapt to the unique non-homophilic patterns of each node, we propose the ProNoG framework for non-homophilic prompt learning, which is equipped with a condition-net to generate a series of prompts conditioned on each node. The node-specific prompts enables fine-grained, node-wise adaptation for the downstream tasks. (3) We perform extensive experiments on ten benchmark datasets, demonstrating the superior performance of ProNoG compared to a suite of state-of-the-art methods.
@@ 2. Related Work
Graph representation learning. GNNs (Kipf and Welling, 2017; Veličković et al., 2018; Fang et al., 2023a; Bo et al., 2023; Fang et al., 2022; Li et al., 2024a, b) are mainstream technique for graph representation learning. They typically operate on a message-passing framework, where nodes iteratively update their representations by aggregating messages received from their neighboring nodes (Xu et al., 2019; Hamilton et al., 2017; Yu et al., 2023; Fang et al., 2023c, b; Jiang et al., 2024a). However, the effectiveness of GNNs heavily relies on abundant task-specific labeled data and requires re-training for various tasks. Inspired by the success of pre-training methods in the language (Dong et al., 2019; Brown et al., 2020; Gao et al., 2021; Schick and Schütze, 2021) and vision (Bao et al., 2022; Zhou et al., 2022b, a; Zang et al., 2022) domains, pre-training methods (Kipf and Welling, 2016; Hu et al., 2020b, a; Lu et al., 2021; Qiu et al., 2020; Velickovic et al., 2019; You et al., 2020; Tan et al., 2024; Mo et al., 2023, 2024; Jiang et al., 2024b) have been widely explored for graphs. These methods first pre-train a graph encoder based on self-supervised tasks, then transfer prior knowledge to downstream tasks. However, all these GNNs and pre-training methods are based on the homophilic assumption, overlooking that real-world graphs are generally non-homophilic.
@@
Non-homophilic graph learning. Many GNNs (Ma et al., 2022; Zhu et al., 2021; Luan et al., 2022; Zhu et al., 2020b; Bo et al., 2021; Zhu et al., 2022; Platonov et al., 2023) have been proposed for non-homophilic graphs, employing methods such as capturing high-frequency signals (Bo et al., 2021), discovering potential neighbors (Pei et al., 2020; Jin et al., 2021), and high-order message passing (Zhu et al., 2020b). Moreover, recent works have explored pre-training on non-homophilic graphs (Xiao et al., 2022; He et al., 2023; Xiao et al., 2023) by capturing neighborhood information to construct unsupervised tasks for pre-training the graph encoder and then transferring prior non-homophilic knowledge to downstream tasks through fine-tuning with task-specific supervision. However, a significant gap exists between the objectives of pre-training and fine-tuning (Liu et al., 2023b; Yu et al., 2024a; Sun et al., 2023b). While pre-training focuses on learning inherent graph attributes without supervision, fine-tuning adapts these insights to downstream tasks based on task-specific supervision. This discrepancy hinders effective knowledge transfer and negatively impacts downstream performance.
@@
Graph prompt learning. Originally developed for the language domain, prompt learning effectively unifies pre-training and downstream objectives (Brown et al., 2020; Liu et al., 2021b; Lester et al., 2021). Recently, graph prompt learning has emerged as a popular alternation to fine-tuning methods (Liu et al., 2023a; Sun et al., 2023a; Tan et al., 2023; Fang et al., 2024; Sun et al., 2022; Yu et al., 2024b, c). These methods first propose a unified template, then design prompts specifically tailored to each downstream task, allowing them to better align with the pre-trained model while keeping the pre-trained parameters frozen. However, current graph prompt learning methods typically assume graphs are homophilic (Yu et al., 2024a; Sun et al., 2023a), neglecting the fact that real-world graphs are generally non-homophilic, exhibiting a mixture of homophilic and heterophilic patterns. Furthermore, these methods usually apply a single prompt for all nodes, overlooking the unique characteristics of each node’s non-homophilic pattern.
@@ 3. Preliminaries
Graph. A graph is defined as $G=(V,E)$, where $V$ represents the set of nodes and $E$ represents the set of edges. The nodes are also associated with a feature matrix $\mathbf{X}\in\mathbb{R}^{|\mathcal{V}|\times d}$, such that $\mathbf{x}_{v}\in\mathbb{R}^{d}$ is a row of $\mathbf{X}$ representing the feature vector for node $v\in V$. For a collection of multiple graphs, we use the notation $\mathcal{G}=\{G_{1},G_{2},\dots,G_{N}\}$.
@@
where $|\mathcal{N}(v)|$ is the set of neighboring nodes of $v$. Note that both $\mathcal{H}(G)$ and $\mathcal{H}(v)$ are in [0,1]. Graphs or nodes with a larger proportion of homophilic edges have a higher homophily ratio.
@@
Problem statement. In this work, we aim to pre-train a graph encoder and develop a prompt learning framework for non-homophilic graphs. More specifically, both the pre-training and prompt learning are not sensitive to the homophilic characteristics of the graph and its nodes.
To evaluate our non-homophilic pre-training and prompt learning, we focus on two common tasks on graph data: node classification and graph classification, in few-shot settings. For node classification within a graph $G=(V,E)$, let $Y$ be the set of node classes. Each node $v_{i}\in V$ has a class label $y_{i}\in Y$. Similarly, for graph classification across a set of graphs $\mathcal{G}$, let $\mathcal{Y}$ be the set of possible graph labels. Each graph $G_{i}\in\mathcal{G}$ has a class label $Y_{i}\in\mathcal{Y}$. In the few-shot setting, there are only $k$ labeled samples per class, where $k$ is a small number (e.g., $k\leq 10$). This scenario is known as $k$-shot classification (Liu et al., 2023a; Yu et al., 2024c, d). Note that the homophily ratio is defined with respect to some predefined set of labels, which may not be related to the class labels in downstream tasks.
@@ 4.1. Theoretical Insights
We focus on contrastive graph pre-training tasks. Given a mainstream contrastive task (You et al., 2020; Qiu et al., 2020; Zhu et al., 2020a; Hassani and Khasahmadi, 2020; Liu et al., 2023a; Yu et al., 2024c), $T=(\{\mathcal{A}_{u}:u\in V\},\{\mathcal{B}_{u}:u\in V\})$, its loss function $\mathcal{L}_{T}$ can be generalized to a standard form below.
@@
where $\mathtt{sim}(\cdot,\cdot)$ represents a similarity function such as cosine similarity (Rahutomo et al., 2012) in our experiment, $\mathcal{A}_{u}$ is the set of positive instances for node $u$, and $\mathcal{B}_{u}$ is the set of negative instances for $u$. The optimization objective of task $T$ in Eq. (4) is to maximize the similarity between $u$ and its positive instances while minimizing the similarity between $u$ and its negative instances. Based on this loss, we further propose the definitions of homophily tasks and homophily samples.
@@
On a graph $G=(V,E)$, consider a triplet $(u,a,b)$ where $u\in V$, $(u,a)\in E$ and $(u,b)\notin E$. The triplet $(u,a,b)$ is a homophily sample if and only $\mathtt{sim}(u,a)>\mathtt{sim}(u,b)$, and it is a non-homophily sample otherwise. ∎
@@
Consider a homophily sample $(u,a,b)$ for some $(u,a)\in E$ and $(u,b)\notin E$, as well as a non-homophily sample $(u,a^{\prime},b^{\prime})$ for some $(u,a^{\prime})\in E$ and $(u,b^{\prime})\notin E$. Let the overall loss with $(u,a,b)$ be $L_{T}$, and that with $(u,a^{\prime},b^{\prime})$ be $L^{\prime}_{T}$. Since $(u,a,b)$ is homophily, we have $\mathtt{sim}(u,a)>\mathtt{sim}(u,b)$, and thus $p(u,a,b)>0.5$. Moreover, since $(u,a^{\prime},b^{\prime})$ is non-homophily, we have $\mathtt{sim}(u,a^{\prime})\leq\mathtt{sim}(u,b^{\prime})$, and thus $p(u,a^{\prime},b^{\prime})\leq 0.5$. Hence, $p(u,a,b)>p(u,a^{\prime},b^{\prime})$, implying that $L_{T}<L^{\prime}_{T}$. ∎
@@
Consider a graph $G=(V,E)$ with a label mapping function $V\rightarrow Y$, and let $y_{v}\in Y$ denote the label mapped to $v\in V$. Suppose the label mapping satisfies that
$\forall u,a,b\in V,y_{u}=y_{a}\land y_{u}\neq y_{b}\Rightarrow\mathtt{sim}(u,a)>\mathtt{sim}(u,b).$
@@
For a homophily task $T=(\{\mathcal{A}_{u}:u\in V\},\{\mathcal{B}_{u}:u\in V\})$, a triplet $(u,a,b)$ for some $u\in V,a\in\mathcal{A}_{u}$ and $b\in\mathcal{B}_{u}$ is a homophily sample with a probability of $P(y_{u}=y_{a})(1-P(y_{u}=y_{b}))$, since $y_{u}=y_{a}\land y_{u}\neq y_{b}$ implies $\mathtt{sim}(u,a)>\mathtt{sim}(u,b)$. Hence, the expected number of homophily samples for $T$ is
@@
For a constant number of nodes with label $y_{u}$, as $\mathcal{H}(G)$ increases, $P(y_{u}=y_{a})$ always increases while $P(y_{u}=y_{b})$ always decreases, leading to a larger $\mathbb{E}_{T}$. ∎

### v6

@@ Abstract.
Graphs are ubiquitous for modeling complex relationships between objects across various fields. Graph neural networks (GNNs) have become a mainstream technique for graph-based applications, but their performance heavily relies on abundant labeled data. To reduce labeling requirement, pre-training and prompt learning has become a popular alternative. However, most existing prompt methods do not distinguish between homophilic and heterophilic characteristics in graphs. In particular, many real-world graphs are non-homophilic—neither strictly nor uniformly homophilic—as they exhibit varying homophilic and heterophilic patterns across graphs and nodes. In this paper, we propose ProNoG, a novel pre-training and prompt learning framework for such non-homophilic graphs. First, we examine existing graph pre-training methods, providing insights into the choice of pre-training tasks. Second, recognizing that each node exhibits unique non-homophilic characteristics, we propose a conditional network to characterize node-specific patterns in downstream tasks. Finally, we thoroughly evaluate and analyze ProNoG through extensive experiments on ten public datasets.
Non-homophilic graph, prompt learning, graph pre-training
Co-first authors. Part of the work was done at the University of Tokyo.
@@
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1; August 3–7, 2025; Toronto, ON, Canada.booktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3690624.3709219isbn: 979-8-4007-1245-6/25/08ccs: Information systems Data miningccs: Computing methodologies Machine learning
@@ 1. Introduction
Graph data are pervasive in real-world applications, such as citation networks, social networks, transportation systems, and molecular graphs. Traditional methods typically train graph neural networks (GNNs) (Kipf and Welling, 2017; Veličković et al., 2018) or graph transformers (Ying et al., 2021; Yun et al., 2019) in a supervised manner. However, they require substantial labeled data and re-training for each specific task.
@@
To mitigate the limitations of supervised methods, pre-training methods have gained significant traction (Hu et al., 2020a; Velickovic et al., 2019; You et al., 2020). They first learn universal, task-independent properties from unlabeled graphs, and then fine-tune the pre-trained models to various downstream tasks using task-specific labels (Velickovic et al., 2019; You et al., 2020). However, a significant gap occurs between the pre-training objectives and downstream tasks, resulting in suboptimal performance (Yu et al., 2024a; Sun et al., 2023b). Moreover, fine-tuning large pre-trained models is costly and still requires sufficient task-specific labels to prevent overfitting. As an alternative to fine-tuning, prompt learning has emerged as a popular parameter-efficient technique for adaptation to downstream tasks (Sun et al., 2022; Liu et al., 2023a; Fang et al., 2024; Yu et al., 2024c). They first utilize a universal template to unify pre-training and downstream tasks. Then, a learnable prompt is employed to modify the input features or hidden embeddings of the pre-trained model to align with the downstream task without updating the pre-trained weights. Since a prompt has far fewer parameters than the pre-trained model, prompt learning can be especially effective in low-resource settings (Yu et al., 2024a).
@@
However, current graph “pre-train, prompt” approaches rely on the homophily assumption or overlook the presence of heterophilic edges. Specifically, the homophily assumption (Ma et al., 2022; Zhu et al., 2020b) states that neighboring nodes should share the same labels, whereas heterophily refers to the opposite scenario where two neighboring nodes have different labels. We observe that real-world graphs are typically non-homophilic, meaning they are neither strictly nor uniformly homophilic and mix both homophilic and heterophilic patterns (Xiao et al., 2022, 2023). In this work, we investigate the pre-training and prompt learning methodology for non-homophilic graphs. We first revisit existing graph pre-training methods, and then propose a Prompt learning framework for Non-homophilic Graphs (or ProNoG in short). The solution is non-trivial, as the notion of homophily encompasses two key aspects, each with its own unique challenge.
@@
First, different graphs exhibit varying degrees of non-homophily. As shown in Fig. 1(a), the Cora citation network is typically considered largely homophilic with 81% homophilic edges [^1] Defined as edges connecting two nodes of the same label; see Eq. (1) in Sect. 3., whereas the Wisconsin web graph links different kinds of webpages, which is highly heterophilic with only 21% homophilic edges. Moreover, the non-homophilic characteristics of a graph also depends on the target label. For example, in a dating network shown in Fig. 1(b), taking gender as the node label, the graph is more heterophilic with $2/7$ homophilic edges. However, taking hobbies as the node label, the graph becomes more homophilic with $4/7$ homophilic edges. Hence, how do we pre-train a graph model irrespective of the graph’s homophily characteristics? In this work, we propose definitions for homophily tasks and homophily samples. We show that pre-training with non-homophily samples increases the loss of any homophily task. Meanwhile, a less homophilic graph results in a higher number of non-homophily samples, subsequently increasing the pre-training loss for homophily tasks. This motivates us to move away from homophily tasks for graph pre-training (Sun et al., 2022; Liu et al., 2023a) and instead choose a non-homophily task (You et al., 2020; Xiao et al., 2022).
@@
Second, different nodes within the same graph are distributed differently in terms of their non-homophilic characteristics. As shown in Fig. 1(c), on each dataset, different nodes within the same graph generally exhibit diverse homophily ratios [^2] Defined as the proportion of a node’s neighbors that share the same label as the node; refer to Eq. (2) in Sect. 3.. Hence, how do we capture the fine-grained, node-specific non-homophilic characteristics? Due to the diverse characteristics across nodes, a one-size-fits-all solution for all nodes would be inadequate. However, existing approaches generally apply a single prompt to all nodes (Sun et al., 2022; Liu et al., 2023a), treating all nodes uniformly. Thus, these methods overlook the fine-grained node-wise non-homophilic characteristics, leading to suboptimal performance. Although some recent works (Tan et al., 2023; Chen et al., 2023) have proposed node-specific prompts, they are not designed to account for the variation in nodes’ non-homophilic charactersitics. Inspired by conditional prompt learning (Zhou et al., 2022a), we propose generating a unique prompt from each node with a conditional network (condition-net) to capture the distinct characteristics of each node. We first capture the non-homophilic patterns of each node by reading out its multi-hop neighborhood. Then, conditioned on these non-homophilic patterns, the condition-net produces a series of prompts, one for each node that reflects its varying non-homophilic characteristics. These prompts can adjust the node embeddings to better align them with the downstream task.
In summary, the contributions of this work are threefold: (1) We observe varying degrees of homophily across graphs, which motivates us to revisit graph pre-training tasks. We provide theoretical insights which guide us to choose non-homophily tasks for graph pre-training. (2) We further observe that, within the same graph, different nodes have diverse distributions of non-homophilic characteristics. To adapt to the unique non-homophilic patterns of each node, we propose the ProNoG framework for non-homophilic prompt learning, which is equipped with a condition-net to generate a series of prompts conditioned on each node. The node-specific prompts enables fine-grained, node-wise adaptation for the downstream tasks. (3) We perform extensive experiments on ten benchmark datasets, demonstrating the superior performance of ProNoG compared to a suite of state-of-the-art methods.
@@ 2. Related Work
In the following, we briefly review the literature on general and non-homophilic graph learning, as well as graph prompt learning.
@@
Graph representation learning. GNNs (Kipf and Welling, 2017; Veličković et al., 2018) are mainstream approaches for graph representation learning. They typically operate on a message-passing framework, where nodes iteratively update their representations by aggregating messages received from their neighboring nodes (Xu et al., 2019; Hamilton et al., 2017; Yu et al., 2023). However, their effectiveness relies on abundant task-specific labeled data and requires re-training for each task. Inspired by the success of the pre-training paradigm in the language (Dong et al., 2019; Brown et al., 2020; Gao et al., 2021; Schick and Schütze, 2021) and vision (Bao et al., 2022; Zhou et al., 2022b, a; Zang et al., 2022) domains, pre-training methods (Kipf and Welling, 2016; Hu et al., 2020b, a; Lu et al., 2021) have been widely explored for graphs. These methods first pre-train a graph encoder based on self-supervised tasks, and subsequently transfer the pre-trained prior knowledge to downstream tasks. However, these pre-training methods often make the homophily assumption, overlooking that real-world graphs are generally non-homophilic.
@@
Non-homophilic graph learning. Many GNNs (Ma et al., 2022; Zhu et al., 2021; Luan et al., 2022) have been proposed for non-homophilic graphs, employing techniques such as capturing high-frequency signals (Bo et al., 2021), discovering potential neighbors (Pei et al., 2020; Jin et al., 2021), and high-order message passing (Zhu et al., 2020b). Moreover, recent works have explored pre-training on non-homophilic graphs (Xiao et al., 2022; He et al., 2023; Xiao et al., 2023) by capturing neighborhood information to construct unsupervised tasks for pre-training a graph encoder, and then transferring prior non-homophilic knowledge to downstream tasks through fine-tuning with task-specific supervision. However, a significant gap exists between the objectives of pre-training and fine-tuning (Liu et al., 2023b; Yu et al., 2024a; Sun et al., 2023b). While pre-training focuses on learning inherent graph properties without supervision, fine-tuning adapts these properties to downstream tasks based on task-specific supervision. This discrepancy hinders effective knowledge transfer and negatively impacts downstream performance.
@@
Graph prompt learning. Originally developed for the language domain, prompt learning effectively unifies pre-training and downstream objectives (Brown et al., 2020; Liu et al., 2021b; Lester et al., 2021). Recently, graph prompt learning has emerged as a popular alternation to fine-tuning methods (Liu et al., 2023a; Sun et al., 2022; Yu et al., 2024b, c). These methods first propose a unified template, then design prompts specifically tailored to each downstream task, allowing them to better align with the pre-trained model while keeping the pre-trained parameters frozen. However, current graph prompt learning methods (Yu et al., 2024a; Sun et al., 2023a) typically do not consider the fact that real-world graphs are generally non-homophilic, exhibiting a mixture of diverse homophilic and heterophilic patterns across nodes. Hence, these methods usually apply a single prompt to all nodes, overlooking the unique non-homophilic pattern of each node.
@@ 3. Preliminaries
Graph. A graph is defined as $G=(V,E)$, where $V$ represents the set of nodes and $E$ represents the set of edges. The nodes are also associated with a feature matrix $\mathbf{X}\in\mathbb{R}^{|V|\times d}$, such that $\mathbf{x}_{v}\in\mathbb{R}^{d}$ is a row of $\mathbf{X}$ representing the feature vector for node $v\in V$. For a collection of multiple graphs, we use the notation $\mathcal{G}=\{G_{1},G_{2},\dots,G_{N}\}$.
@@
where $\mathcal{N}(v)$ is the set of neighboring nodes of $v$. Note that both $\mathcal{H}(G)$ and $\mathcal{H}(v)$ fall in $[0,1]$. Graphs or nodes with a larger proportion of homophilic edges have a higher homophily ratio.
@@
Problem statement. In this work, we aim to pre-train a graph encoder and develop a prompt learning framework for non-homophilic graphs. More specifically, both the pre-training and prompt learning are not sensitive to the diverse non-homophilic characteristics of the graph and its nodes.
To evaluate our non-homophilic pre-training and prompt learning, we focus on two common tasks on graph data: node classification and graph classification, in few-shot settings. For node classification within a graph $G=(V,E)$ over a set of node classes $Y$, each node $v_{i}\in V$ has a class label $y_{i}\in Y$. Similarly, for graph classification across a graph collection $\mathcal{G}$ with class labels $Y$, each graph $G_{i}\in\mathcal{G}$ has a class label $Y_{i}\in Y$. In the few-shot setting, there are only $k$ labeled samples per class, where $k$ is a small number (e.g., $k\leq 10$). This scenario is known as $k$-shot classification (Liu et al., 2023a; Yu et al., 2024c, d). Note that the homophily ratio is defined with respect to some predefined set of labels, which may or may not be related to the class labels in downstream tasks.
@@ 4.1. Theoretical Insights
We focus on contrastive graph pre-training tasks. Consider a mainstream contrastive task (You et al., 2020; Qiu et al., 2020; Zhu et al., 2020a; Hassani and Khasahmadi, 2020; Liu et al., 2023a; Yu et al., 2024c), $T=(\{\mathcal{A}_{u}:u\in V\},\{\mathcal{B}_{u}:u\in V\})$, where $\mathcal{A}_{u}$ is the set of positive instances for node $u$, and $\mathcal{B}_{u}$ is the set of negative instances for $u$. Its loss function $\mathcal{L}_{T}$ can be standardized (Yu et al., 2024c) to a similar form as follows.
@@
where $\mathtt{sim}(\cdot,\cdot)$ represents a similarity function such as cosine similarity in our experiments. The optimization objective of task $T$ in Eq. (4) is to maximize the similarity between $u$ and its positive instances while minimizing the similarity between $u$ and its negative instances. Based on this loss, we further propose the definitions of homophily tasks and homophily samples.
@@
On a graph $G=(V,E)$, consider a triplet $(u,a,b)$ where $u\in V$, $(u,a)\in E$ and $(u,b)\notin E$. The triplet $(u,a,b)$ is a homophily sample if and only if $\mathtt{sim}(\mathbf{h}_{u},\mathbf{h}_{a})>\mathtt{sim}(\mathbf{h}_{u},\mathbf{h}_{b})$, and it is a non-homophily sample otherwise. ∎
@@
Consider a homophily sample $(u,a,b)$ for some $(u,a)\in E$ and $(u,b)\notin E$, as well as a non-homophily sample $(u,a^{\prime},b^{\prime})$ for some $(u,a^{\prime})\in E$ and $(u,b^{\prime})\notin E$. Let the overall loss with $(u,a,b)$ be $L_{T}$, and that with $(u,a^{\prime},b^{\prime})$ be $L^{\prime}_{T}$. Since $(u,a,b)$ is a homophily sample, we have $\mathtt{sim}(\mathbf{h}_{u},\mathbf{h}_{a})>\mathtt{sim}(\mathbf{h}_{u},\mathbf{h}_{b})$, and thus
$p(u,a,b)=\textstyle\frac{\mathtt{sim}(\mathbf{h}_{u},\mathbf{h}_{a})}{\mathtt{sim}(\mathbf{h}_{u},\mathbf{h}_{a})+\mathtt{sim}(\mathbf{h}_{u},\mathbf{h}_{b})}>0.5.$
Moreover, since $(u,a^{\prime},b^{\prime})$ is a non-homophily sample, we have $\mathtt{sim}(\mathbf{h}_{u},\mathbf{h}_{a}^{\prime})\leq\mathtt{sim}(\mathbf{h}_{u},\mathbf{h}_{b}^{\prime})$, and thus $p(u,a^{\prime},b^{\prime})\leq 0.5$. Hence, $p(u,a,b)>p(u,a^{\prime},b^{\prime})$, implying that $L_{T}<L^{\prime}_{T}$. ∎
@@
Consider a graph $G=(V,E)$ with a label mapping function $V\rightarrow Y$, where $y_{v}\in Y$ is the label mapped to $v\in V$. Suppose that the label mapping and node similarity are consistent, i.e.,
$\forall u,a,b\in V,y_{u}=y_{a}\land y_{u}\neq y_{b}\Rightarrow\mathtt{sim}(\mathbf{h}_{u},\mathbf{h}_{a})>\mathtt{sim}(\mathbf{h}_{u},\mathbf{h}_{b}).$
@@
For a homophily task $T=(\{\mathcal{A}_{u}:u\in V\},\{\mathcal{B}_{u}:u\in V\})$, a triplet $(u,a,b)$ for some $u\in V,a\in\mathcal{A}_{u}$ and $b\in\mathcal{B}_{u}$ is a homophily sample with a probability of $P(y_{u}=y_{a})(1-P(y_{u}=y_{b}))$, since $y_{u}=y_{a}\land y_{u}\neq y_{b}$ implies $\mathtt{sim}(\mathbf{h}_{u},\mathbf{h}_{a})>\mathtt{sim}(\mathbf{h}_{u},\mathbf{h}_{b})$. Hence, the expected number of homophily samples for $T$ is
@@
For a constant number of nodes with label $y_{u}$, as $\mathcal{H}(G)$ increases, $P(y_{u}=y_{a})$ increases while $P(y_{u}=y_{b})$ decreases, leading to a larger $\mathbb{E}_{T}$. ∎